### v2

@@ Abstract
Organic electro-optic materials have demonstrated promising performance in developing electro-optic phase shifters. Their integration with other silicon photonic processes, nanofabrication complexities, and durability remain to be developed. While the required poling step in electro-optic polymers limits their potential and utilization on a large scale, devices made of paraelectric nematic liquid crystals suffer from slow bandwidth. In ferroelectric nematic liquid crystals, we report an additional GHz-fast phase shift that ultimately allows for significant second-order nonlinear optical coefficients and the related Pockels effect. It avoids poling issues and can pave the way for hybrid silicon-organic systems with CMOS-foundry compatibility. We report DC and AC modulation efficiencies of 0.25 Vmm and 25.7 Vmm, respectively, an on-chip insertion loss of 2.1 dB, and an electro-optic bandwidth of >4.18 GHz, employing improved light-matter interaction in a waveguide architecture that calls for only one lithography step.
@@
Silicon and silicon nitride photonics have enabled several recent technological advancements, and future roadmaps have been designed based on their capabilities and integration with established CMOS foundries [1]. The electro-optic phase shifter (EOPS) [2] is one of the most fundamental components upon which other essential building blocks in photonic integrated circuits (PIC) are based, including electro-optic (EO) switches [3], EO modulators [4], optical interconnects [5], and comb generators [6]. Traditional phase shift mechanisms often show poor overall performance, such as limited EO bandwidth (BW) and high static energy consumption () when leveraging the thermo-optic effect or free-carrier dispersion effect with traveling wave modulators. Free carrier dispersion also leads to doping-induced insertion loss (IL). Innovations in hybrid material platforms promise viable solutions to traditional Si photonics (SiP) shortcomings. However, they typically face significant challenges, such as several complex nanofabrication stages, scaling, and interaction with current CMOS foundries. Modulators based on indium tin oxide (ITO) have recently experienced notable improvements [4,7-9]. The additional operations necessary for ITO/insulator thin-film deposition, demanding quality control methods, and their low extinction ratio (ER) suggest further R&D needs [10]. Carrier-free phase shift processes, such as the Pockels effect, are among the more appealing techniques, especially regarding EO bandwidth and energy consumption. These metrics and low IL can excel in LiNbO3 on silicon (LNOI) [11-12]. Due to the low EO coefficient of LiNbO3, however, phase shifters based on LNOI require either a large footprint or a high drive voltage (), where is the voltage necessary to produce shift. A material system based on barium titanate (BTO) can simultaneously meet the requirements of low IL and high EO bandwidth [13-14]. Despite having a significant EO coefficient, BTO is unable to fully utilize the improved light-matter interactions (LMI) provided by some waveguide topologies, such as slot waveguides, failing to give the record-low modulation efficiencies (), where is the EOPS effective length [15]. BTO also involves more meticulous fabrication steps, loses its high EO efficiency at cryogenic temperatures, and necessitates process controls for optimum domain development [16].
@@
Furthermore, the preceding approaches cannot meet the trade-offs between various device metrics of a particular modulator, such as footprint, EO and optical BW, IL, , , and dynamic energy consumption () [17-18]. Hybrid integration of Si with organic electro-optic (OEO) materials has been proposed to overcome these shortcomings. They have advantages over their inorganic equivalents due to a combination of their EO and physicochemical characteristics. The EO properties of OEO materials, including the Pockels factor () and optical loss, have continuously progressed over the last two decades ( as large as 9000 pmV-1 and absorption coefficients as low as have been reported [19], where is the OEO refractive index and represents the Pockels coefficient). Thanks to their liquid state, an OEO can infiltrate nanostructures during the coating step and benefit from the enhanced LMI. Along with the resultant high modulation efficiencies, the low dielectric dispersion [20] of OEO promises better modulators. As a result, <1 V, sub-mm-long Mach-Zehnder modulators (MZM) with IL < 1 dB, EO bandwidth surpassing 100 GHz, and 100 Jbit-1 have been enabled [21]. Such an all-inclusive, miniaturized modulator can work in extreme temperatures (as high as 383 K (110C) [22] and as low as 4 K [23,24]), which encompass the crucial requirements for a state-of-the-art modulator.
@@
While EO polymers have attracted increased attention mainly because of their remarkable EO characteristics, their application to photonic-integrated circuits (PIC) has challenges. Multiple process steps and often hazardous solvents are needed to prepare EO polymer thin films, necessitating equipment for precise mixing, coating, and drying. In addition, to align chromophores in an EO polymer material, an electro-thermal poling step is necessary. On a larger scale, dedicated DC lines may be required depending on the waveguide geometry and accompanying LMI. Despite considerable breakthroughs in producing thermally and temporally stable polymers, not all poled polymers are resilient to high temperatures ([22]), aging ([25]), and poling-induced optical loss ([26]). For practical applications, the devices may still require a certain level of hermetic sealing to become temperature- and humidity-resistant. EO polymers are also recognized for their lower resistance at higher temperatures, which coincides with detrimental leakage currents. It leads to decreased EO efficiency, which has a direct impact on device performance. Additional nanofabrication steps have been identified as advantageous to resolving this issue, such as exploiting a thin charge-barrier layer comprised of alumina or titania created by atomic layer deposition, [27,28]. Poling also limits high-temperature excursions typical of back-end-of-line processing since such excursions would destroy the poling-induced order. The problems above could increase the overall cost, add complexity, and jeopardize the integration of polymer-based hybrid platforms.
@@
We report the proofs of a considerably faster EO effect in FN-LC, utilizing the nanoscale properties of our SOH structure. We specifically show that FN-LC can be integrated with similar ease as PN-LCs onto a SiP platform, that alignment of the polar axis can be easily achieved over the length of the device, and that a significant EO response can be observed. This paper proposes a hybrid material platform for PICs comprising numerous moderate-speed EOPS that prioritize large-scale scaling, integrability, and ease of fabrication. This includes non-modulation applications, such as switches [3,39], transducers [40], and weight banks [41] for neuromorphic PICs [42] and quantum computing [43].
@@ Results
Waveguides such as Si [44] and plasmonic [45] slotted, sub-wavelength grating (SWG) [46], photonic-crystal [47], and thinned [22], are traditionally employed as EOPS in modulators. However, they often add excessive loss [48], limit optical BW [48], suffer from poor infiltration, and poling [27,49]. Of particular interest for the current work is the extra fabrication stages slotted-based devices demand [40], some of which are not quite CMOS-compatible. For instance, opening oxide cladding over isolated Si islands of SWG or metallic slots of plasmonic waveguides may violate design rules in wafer nanofabrication. Also, exploiting the benefits of structures such as thinned waveguides necessitates taper converters from complete to partial etch Si [22]. Geometries required for efficient and low-loss mode converters, such as strip-to-slot, are not always accessible through nanofabrication runs or have integration challenges with the rest of the PIC. On the other hand, a non-slotted waveguide does not have these issues at the cost of offering a weak LMI [40], thus requiring a high voltage for poling and driving. The device size and EO bandwidth must then be compromised to keep the voltage within the ranges supplied by CMOS drivers.
We propose a finger-loaded non-slotted (FNS) waveguide consisting of two strip-to-finger mode converters to resolve this trade-off, as seen in Fig. 1(a). In analogy to segmented slotted [50] and non-slotted [51] topologies, this structure eliminates the misalignment problem seen in strip-loaded slot waveguides by offering at least one fewer step for lithography and etching. The periodicity and duty cycle of the subwavelength structure of the FNS waveguide are restricted by fabrication capabilities yet selected to produce a better balance between electrical and optical loss penalties, another prevalent difficulty in slot waveguides. The waveguide core, sandwiched between fingers, is narrowed to 300 nm to localize light in the gap between them by a factor of , as seen in Fig. 2.
Other than the so-called optical field confinement factor (), what matters more is the overlap integral () between optical () and electrical () modal fields. Considering the high , the short effective distance between the two electrodes () brought by the fingers adjacent to the core, the FNS waveguide offers competitive modulation efficiency-loss product (i.e., ) compared to non-slotted waveguide (see Methods). Besides, our simulations revealed that the 10 -long strip-to-finger mode converter does not substantially impact total EOPS loss (conversion efficiency 99.86%). Thanks to its porous structure, FNS topology makes the most of the capillary force action to mitigate material infiltration, another ubiquitous issue in slot waveguides that frequently calls for oxide undercutting [52].
@@
The constructed devices must then be aligned with an electric field. Because of its pre-ordered molecular orientation and sensitivity to an external field, FN-LC was found to be fully aligned with > 4 times smaller than required for poling EO polymers [53], in similar structures. The process was carried out at room temperature; no precise temperature monitor was necessary. These relaxed voltage and temperature conditions have additional advantages. First, an inert gas such as N2 is not required around the chip during alignment to avoid oxidation-driven degradation. The FN-LC alignment process is relatively safer because there is no off-gassing. Within the scope of our characterization setup, the device’s performance exhibited little temperature susceptibility thanks to the non-electro-thermal alignment technique. To be more precise, no performance aging was discovered for the FN-LC-coated devices over 150 hours of continuous operation in the range of 10C under full optical and electrical load. EO polymers lose their poling efficiency beyond their glass transition temperature (C). To show that FN-LC is useful at these temperatures, more accelerated aging tests must be done in the future. The lower voltage required for alignment simplifies design concerns for the large-scale integration of FN-LC-enabled photonic devices in which the same RF lines can be utilized.
@@
Monitoring optical loss and leakage current throughout the alignment process provides critical information. As seen in Fig. 3(a)-(b), changes in optical loss followed by saturation indicate that domain formation has been completed. In some cases, the change was permanent, indicating that the optic axis continues to remain in a vertical posture. As shown in Fig. 3 (b), the fully-aligned FN-LC MZM showed 2 to 3 dB less material absorption loss compared to the material with unaligned dipoles, which we attribute to the formation of a monodomain over the length of the device and reduced scattering loss (see Methods). The alignment-induced improvement in IL contradicts that seen commonly in EO polymers.
Following the alignment, we performed DC characterization of the optical transmission, as shown in Fig. 3(c), achieving an optical ER of 26 dB and 0.5 V. Various effects, including the birefringence effect, are responsible for this considerable shift [34-35,38,54]. The optical transmission of the proposed MZM shown in Fig. 3(d) reads an IL of 2.1 dB, which we attribute mostly to the propagation loss through the non-optimized subwavelength structure of the deployed FNS waveguides. We also found that FN-LC has approximately the same absorption loss as an available EO polymer at the wavelength of operation.
@@
A more comprehensive figure of merit for a given EOPS is the pure phase efficiency (). The metric is defined as ), where and are the real and imaginary parts of the complex refractive index (), respectively. and loss can be correlated to the estimated values for and . The optical ER, , comprise a comprehensive assessment of EOPS appropriateness. Table LABEL:FOMtable summarizes the abovementioned figure of merits for an EOPS built on hybrid material platforms, including FN-LC and the conventional pn-junction technology. For our proposed FNS waveguide, significantly higher records up to 88.9 have been achieved. We believe the proposed EOPS’ improved phase shift performance stems from combining our device design and the material.
@@
With a high-speed phase shift mechanism, FN-LC might demonstrate significant benefits over traditional PN-LC, in which the slower mechanisms mostly dominate the response. Various aspects of the frequency response characterization of the FN-LC-coated devices are shown in Fig. 4, and the measurement details are covered in Methods. SWG-based grating couplers (GC) were used for light injection into the devices for this step [55-57]. The S-parameter collected by a Vector Network Analyzer (VNA) indicates that 4.18 GHz. This achievement demonstrates a phase shift mechanism with a rise/fall time in the nanosecond range, while PN-LC is limited to functioning in the millisecond range [2]. Accordingly, we employed data and an analytical approach (see Methods) to evaluate the EO properties of the FN-LC material and the proposed EOPS. As a result, an EO coefficient of 24 pmV-1 and an AC modulation efficiency of 25.7 (Vmm) were estimated. The metrics are still significantly lower than the maximum record for EO polymers. This, being the first demonstration, promises further improvement in the future.
@@
Considering the sub-mm phase shifters we used in each arm of the MZM, the mismatch between the optical and RF effective indices is not a limiting factor for the EO bandwidth [18]. The EO bandwidth is primarily limited by the weakly-doped nature of fingers in the fabricated FNS waveguides. As shown by Fig. 1, an RC time constant of GHz is anticipated, which agrees with Fig. 4(d). As an immediate remedy, one could reduce finger lengths (i.e., in Fig. 1) or place electrodes closer to the finger’s end. As a secondary factor, the electrical wire bonds (EWB) that apply the RF signal could also have deteriorated the EO bandwidth. Unlike in slot waveguides, where some optical mode is confined in the two Si rails, less modal field propagates inside fingers in the FNS architecture. A single degree of ion implantation would consequently be sufficient to alleviate the poor shown in this work while maintaining the RF-optical loss trade-off. Besides, as shown in Fig. 2(d), the FNS waveguide’s twin slots reduce the overall . That means, less doping can alternatively achieve a desired , which instead would help improving the device’s (Supplementary Note 3).
@@
We discovered that producing dependable EWBs with less complexity was relatively more straightforward than EO polymer-coated samples. The FN-LC phase shift mechanism appears to swing from a strong birefringence-driven effect at low frequencies to a weaker but faster Pockels effect at higher frequencies. Complex phase manipulation is usually required to develop advanced modulation systems, such as coherent transceivers [58]. Inefficient thermal heaters are frequently used to produce a slow phase shift, as illustrated in Fig. 4(c). Including slow and fast EOPS in a single material could be an appealing feature of the FN-LC-based platform.
@@
We evaluated the potential large-scale capability of our EOPS structures by creating a SiP chip with 108 MZMs on a 0.5 mm2 area, as depicted in Fig. 5. A subset of devices was photonically wire-bonded (PWB) to an external laser source via a 16-channel fiber array (FA) block for demonstration purposes. The PWB routes were established using adiabatic surface tapers at the chip edge, followed by selective patterning of FN-LC on active components in a post-processing step. Because of its high viscosity, FN-LC can be applied locally, allowing for more controlled integration with the rest of the chip’s components. Thanks to the reduced alignment voltage enabled by utilizing FNS waveguides in our Si-FN-LC design, the step can be paired with a , which is used to bias modulators at their quadrature point. Common EWB and a printed circuit board (PCB) were used to conduct alignment and drive the modulators. The demonstrated FN-LC SiP chip can be viewed as a prototype towards large-scale integration of organic material-enabled PICs, which we credit partly to overcoming the poling difficulties by utilizing the poling-free materials described here. A hypothetical EO polymeric version of this chip necessitates on-chip routing considerations and a separate PCB with potentially higher-voltage handling capabilities for poling.
Towards a truly integrated SiP-FN-LC chip on a large scale, one would require to show the cooperation of several EOPs in relevant scenarios such as the one reported in [59], which require a comprehensive electrical packaging scheme and will be covered in future work. Previously, we have also shown successful integration of other photonic components such as distributed feedback laser [60] and semiconductor laser amplifier [61] using similar photonic packaging technique shown in this work. Furthermore, we anticipate some sealing would be beneficial to accomplishing a properly packaged FN-LC SiP chip despite our early observations of the insensitivity of FN-LC to agents such as humidity and oxygen.

### v3

@@ Abstract
Organic electro-optic materials have demonstrated promising performance in developing electro-optic phase shifters. Their integration with other silicon photonic processes, nanofabrication complexities, and durability remains to be developed. While the required poling step in electro-optic polymers limits their potential and utilization on a large scale, devices made of paraelectric nematic liquid crystals suffer from slow bandwidth. In ferroelectric nematic liquid crystals, we report an additional GHz-fast phase shift that ultimately allows for significant second-order nonlinear optical coefficients related to the Pockels effect. It avoids poling issues and can pave the way for hybrid silicon-organic systems with CMOS-foundry compatibility. We report DC and AC modulation efficiencies of 0.25 Vmm (from liquid crystal orientation) and 25.7 Vmm (from Pockels effect), respectively, an on-chip insertion loss of 2.6 dB, and an electro-optic bandwidth of >4.18 GHz, employing improved light-matter interaction in a waveguide architecture that calls for only one lithography step.
@@
Silicon and silicon nitride photonics have enabled several recent technological advancements, and future roadmaps have been designed based on their capabilities and integration with established CMOS foundries [1]. The electro-optic phase shifter (EOPS) [2] is one of the most fundamental components upon which other essential building blocks in photonic integrated circuits (PIC) are based, including electro-optic (EO) switches [3], EO modulators [4], optical interconnects [5], and comb generators [6]. Traditional phase shift mechanisms often show poor overall performance, such as limited EO bandwidth (BW) and high static power consumption () when leveraging the thermo-optic effect or free-carrier dispersion effect with travelling wave modulators. Free carrier dispersion also leads to doping-induced insertion loss (IL). Innovations in hybrid material platforms promise viable solutions to traditional Si photonics (SiP) shortcomings. However, they typically face significant challenges, such as several complex nanofabrication stages, scaling, and interaction with current CMOS foundries. Modulators based on indium tin oxide (ITO) have recently experienced notable improvements [7]. The additional operations necessary for ITO/insulator thin-film deposition, demanding quality control methods, and their low extinction ratio (ER) suggest further R&D needs [8]. Carrier-free phase shift processes, such as the Pockels effect, are among the more appealing techniques, especially regarding EO bandwidth and energy consumption. These metrics and low IL can excel in LiNbO3 on silicon (LNOI) [9-10]. Due to the low EO coefficient of LiNbO3, however, phase shifters based on LNOI require either a large footprint or a high drive voltage (), where is the voltage necessary to produce shift. A material system based on barium titanate (BTO) can simultaneously meet the requirements of low IL and high EO bandwidth [11-12]. Despite having a significant EO coefficient, BTO is unable to fully utilize the improved light-matter interactions (LMI) provided by some waveguide topologies, such as slot waveguides, failing to give the record-low modulation efficiencies (), where is the EOPS effective length [13]. BTO also involves meticulous fabrication steps, loses its high EO efficiency at cryogenic temperatures, and necessitates process controls for optimum domain development [14].
@@
Furthermore, the preceding approaches cannot meet the trade-offs between various device metrics of a particular modulator, such as footprint, EO and optical BW, IL, , and dynamic energy consumption () [15-16]. Hybrid integration of Si with organic electro-optic (OEO) materials has been proposed to overcome these shortcomings. They have advantages over their inorganic equivalents due to a combination of their EO and physicochemical characteristics. The EO properties of OEO materials, including the Pockels factor () and optical loss, have continuously progressed over the last two decades ( as large as 9000 pmV-1 and absorption coefficients as low as have been reported [17], where is the OEO refractive index and represents the Pockels coefficient). Thanks to their liquid state, an OEO can infiltrate nanostructures during the coating step and benefit from the enhanced LMI. OEO’s low dielectric dispersion [18] promises better modulators and the resultant high modulation efficiencies. As a result, <1 V, sub-mm-long Mach-Zehnder modulators (MZM) with IL < 1 dB, EO bandwidth surpassing 100 GHz, and 100 Jbit-1 have been enabled [19]. Such an all-inclusive, miniaturized modulator can work in extreme temperatures (as high as 383 K (110C) [20] and as low as 4 K [21,22]), which encompass the crucial requirements for a state-of-the-art modulator.
@@
While EO polymers have attracted increased attention mainly because of their remarkable EO characteristics, their application to photonic-integrated circuits (PIC) has challenges. Multiple process steps and often hazardous solvents are needed to prepare EO polymer thin films, necessitating equipment for precise mixing, coating, and drying. In addition, to align chromophores in an EO polymer material, an electro-thermal poling step is necessary. On a larger scale, dedicated DC lines may be required depending on the waveguide geometry and accompanying LMI. Despite considerable breakthroughs in producing thermally and temporally stable polymers, not all poled polymers are resilient to high temperatures ([20]), aging ([23]), and poling-induced optical loss ([24]). For practical applications, the devices may still require a certain level of hermetic sealing to become temperature- and humidity-resistant. EO polymers are also recognized for their lower resistance at higher temperatures, which coincides with detrimental leakage currents. It leads to decreased EO efficiency, which has a direct impact on device performance. Additional nanofabrication steps have been identified as advantageous to resolving this issue, such as exploiting a thin charge-barrier layer of alumina or titania created by atomic layer deposition [25,26]. Poling also limits high-temperature excursions typical of back-end-of-line processing since such excursions would destroy the poling-induced order. The problems above could increase the overall cost, add complexity, and jeopardize the integration of polymer-based hybrid platforms.
@@
We report demonstrating a considerably faster EO effect in FN-LC, utilizing the nanoscale properties of our SOH structure. We specifically show that FN-LC can be integrated with similar ease as PN-LCs onto a SiP platform, that alignment of the polar axis can be easily achieved over the length of the device, and that a significant EO response can be observed. This paper proposes a hybrid material platform for PICs comprising numerous moderate-speed EOPS that prioritize large-scale scaling, integrability, and ease of fabrication. It includes non-modulation applications, such as switches [3,37], transducers [38], and weight banks [39] for neuromorphic PICs [40] and quantum computing [41].
@@ Results
Waveguides such as Si [42] and plasmonic [43] slotted, sub-wavelength grating (SWG) [44], photonic-crystal [45], and thinned [20], are traditionally employed as EOPS in modulators. However, they often add excessive loss [46], limit optical BW [46], and suffer from poor infiltration and poling [25,47]. Of particular interest for the current work is the extra fabrication stages slotted-based devices demand [38], some of which are not quite CMOS-compatible. For instance, opening oxide cladding over isolated Si islands of SWG or metallic slots of plasmonic waveguides may violate design rules in wafer nanofabrication. Also, exploiting the benefits of structures such as thinned waveguides necessitates taper converters from complete to partial etch Si [20]. Geometries required for efficient and low-loss mode converters, such as strip-to-slot, are not always accessible through nanofabrication runs or have integration challenges with the rest of the PIC. On the other hand, a non-slotted waveguide does not have these issues at the cost of offering a weak LMI [38], thus requiring a high voltage for poling and driving. The device size and EO bandwidth must then be compromised to keep the voltage within the ranges supplied by CMOS drivers.
We propose a finger-loaded strip (FLS) waveguide consisting of two strip-to-finger mode converters to resolve this trade-off, as seen in Fig. 1(a). In analogy to segmented slotted [48] and non-slotted [49] topologies, this structure eliminates the misalignment problem seen in strip-loaded slot waveguides by offering at least one fewer step for lithography and etching. The periodicity and duty cycle of the subwavelength structure of the FLS waveguide is restricted by fabrication capabilities yet selected to produce a better balance between electrical and optical loss penalties, another prevalent difficulty in slot waveguides. The waveguide core, sandwiched between fingers, is narrowed to 300 nm. The optical field confinement factor () of the localized light in the dual gaps between fingertips and the core is shown in Fig. 2(d), and the simulated values are provided in Methods.
Other than the so-called optical field confinement factor , what determines the modulation sensitivity of a device based on the waveguide is the overlap integral () between optical () and electrical () modal fields. Considering the high , the short effective distance between the two electrodes () brought by the fingers adjacent to the core, the FLS waveguide offers competitive modulation efficiency-loss product (i.e., ) compared to a non-slotted waveguide (see Methods). Besides, our simulations revealed that the 10 -long strip-to-finger mode converter does not substantially impact total EOPS loss (conversion efficiency ). Thanks to its porous structure, FLS topology makes the most of the capillary force action to mitigate material infiltration, another ubiquitous issue in slot waveguides that frequently calls for oxide undercutting [50].
@@
The constructed devices must then be aligned with an electric field. Because of its pre-ordered molecular orientation and sensitivity to an external field, FN-LC was found to be fully aligned with times smaller than required for poling EO polymers [51], in similar structures. The process was carried out at room temperature; no precise temperature monitor was necessary. These relaxed voltage and temperature conditions have additional advantages. First, an inert gas such as N2 is not required around the chip during alignment to avoid oxidation-driven degradation. The FN-LC alignment process is relatively safer because there is no off-gassing. Within the scope of our characterization setup, the device’s performance exhibited little temperature susceptibility thanks to the non-electro-thermal alignment technique. To be more precise, no performance aging was discovered for the FN-LC-coated devices over 150 hours of continuous operation in the range of 10C under full optical and electrical load. EO polymers lose their poling efficiency beyond their glass transition temperature (C). To show that FN-LC is helpful at these temperatures, more accelerated aging tests must be done in the future. The lower voltage required for alignment simplifies design concerns for the large-scale integration of FN-LC-enabled photonic devices in which the same RF lines can be utilized.
@@
Monitoring optical loss and leakage current throughout the alignment process provides critical information. As seen in Fig. 3(a)-(b), changes in optical loss followed by saturation indicate that domain formation has been completed. In some cases, the change was permanent, indicating that the optic axis continues to remain in a vertical posture. As shown in Fig. 3 (b), the fully-aligned FN-LC MZM showed 2 to 3 dB less material absorption loss compared to the material with unaligned dipoles, which we attribute to the formation of a monodomain over the length of the device and reduced scattering loss (see Methods). The alignment-induced improvement in IL is opposite to that seen commonly in EO polymers, where the loss increases with voltage.
Following the alignment, we performed DC characterization of the optical transmission, as shown in Fig. 3(c), achieving an optical ER of 26 dB and 0.5 V. Various processes cause this significant shift, the most prominent of which is the birefringence effect. The optical transmission of the proposed MZM, shown in Fig. 3(d), demonstrates an IL of dB. Based on passive measurements of similar structures coated with a low-loss polymer featuring approximately the same index of refraction, we found dB of the IL primarily attributed to the propagation loss through the non-optimized subwavelength structure of the deployed FLS waveguides and the rest dB stemmed from material absorption at the wavelength of operation. Additionally, the birefringence effect contributes to an increase of around 1 dB, as indicated in Fig.3(d). Nonetheless, just a is required during RF operation, implying an additional loss of dB, resulting in a total IL of dB. We also found that FN-LC has approximately the same absorption loss as an available EO polymer at the wavelength of operation. Aside from fabrication nonidealities, a physical path length difference of integrated into one of the arms has led to a free spectral range (FSR) of , as illustrated by Fig. 3(e). It agrees with the theoretical FSR of nm, where is the group index of the FLS waveguide.
@@
Pure phase efficiency is a seldom reported but very important figure of merit for EOPS for communications applications (). The metric is defined as ), where and are the real and imaginary parts of the complex refractive index (), respectively. and loss can be correlated to the estimated values for and . The optical ER, , comprise a comprehensive assessment of EOPS appropriateness. Table LABEL:FOMtable summarizes the abovementioned figure of merits for an EOPS built on hybrid material platforms, including FN-LC and the conventional pn-junction technology. For our proposed FLS waveguide, significantly higher records up to 88.9 have been achieved, which we credit largely to the material properties.
@@
With a high-speed phase shift mechanism, FN-LC might demonstrate significant benefits over traditional PN-LC, in which the slower mechanisms mostly dominate the response. Various aspects of the frequency response characterization of the FN-LC-coated devices are shown in Fig. 4, and the measurement details are covered in Methods. This step used SWG-based grating couplers (GC) for light injection into the devices [53-55]. The S-parameter collected by a Vector Network Analyzer (VNA) indicates that a -6dB EO bandwidth 4.18 GHz. This achievement demonstrates a phase shift mechanism with a rise/fall time in the nanosecond range, while PN-LC is limited to functioning in the millisecond range [2]. Accordingly, we employed data and an analytical approach (see Methods) to evaluate the EO properties of the FN-LC material and the proposed EOPS. As a result, an EO coefficient of 24 pmV-1 and an AC modulation efficiency of 25.7 (Vmm) were estimated. The metrics are still significantly lower than the maximum record for EO polymers. This, being the first demonstration, promises further improvement in the future.
@@
Considering the sub-mm phase shifters we used in each arm of the MZM, the mismatch between the optical and RF effective indices is not a limiting factor for the EO bandwidth [16]. The EO bandwidth is primarily limited by the weakly doped nature of fingers in the fabricated FLS waveguides. As shown by Fig. 1, a -3dB EO bandwidth () limited by the RC time constant () of GHz is anticipated, which agrees with Fig. 4(d). As an immediate remedy, one could reduce finger lengths (i.e., in Fig. 1) or place electrodes closer to the finger’s end. As a secondary factor, the electrical wire bonds (EWB) that apply the RF signal could also have deteriorated the EO bandwidth. Unlike in slot waveguides, where some optical mode is confined in the two Si rails, less modal field propagates inside fingers in the FLS architecture. A single ion implantation step would consequently be sufficient to alleviate the poor shown in this work while maintaining the RF-optical loss trade-off. Figure 4(d) illustrates superimposed results from applying FN-LC to a properly doped slotted waveguide to validate this hypothesis (data extracted from [67]). Besides, as shown in Fig. 2(d), the FLS waveguide’s twin slots reduce the overall . That means, less doping can alternatively achieve a desired , which instead would help improving the device’s (Supplementary Note 3).
@@
We discovered that producing dependable EWBs with less complexity was relatively more straightforward than EO polymer-coated samples. The FN-LC phase shift mechanism appears to swing from a strong birefringence-driven effect at low frequencies to a weaker but faster Pockels effect at higher frequencies. Complex phase manipulation is usually required to develop advanced modulation systems, such as coherent transceivers [56]. Inefficient thermal heaters are frequently used to produce a slow phase shift (e.g., ), as illustrated in Fig. 4(c). Including slow and fast EOPS in a single material could be an appealing feature of the FN-LC-based platform.
@@
We evaluated the potential large-scale capability of our EOPS structures by creating a SiP chip with 108 MZMs on a 0.5 mm2 area, as depicted in Fig. 5. A subset of devices was photonically wire-bonded (PWB) to an external laser source via a 16-channel fiber array (FA) block for demonstration purposes. The PWB routes were established using adiabatic surface tapers at the chip edge, followed by selective patterning of FN-LC on active components in a post-processing step. Because of its high viscosity, FN-LC can be applied locally, allowing for more controlled integration with the rest of the chip’s components. Thanks to the reduced alignment voltage enabled by utilizing FLS waveguides in our Si-FN-LC design, the step can be paired with a , which is used to bias modulators at their quadrature point. Common EWB and a printed circuit board (PCB) were used to conduct alignment and drive the modulators. The demonstrated FN-LC SiP chip can be viewed as a prototype towards large-scale integration of organic material-enabled PICs, which we credit partly to overcoming the poling difficulties by utilizing the poling-free materials described here. A hypothetical EO polymeric version of this chip necessitates on-chip routing considerations and a separate PCB with potentially higher-voltage handling capabilities for poling.
The dual-phase shift of FN-LC provides a significant benefit for large-scale integration of dense EOPSs, with an estimated W required to maintain perfect alignment and drive the MZM at the quadrature point. At the frequencies of interest in this work, the proposed MZM’s comparatively short arm length of mm eliminates the necessity for doubly-terminated travelling wave electrodes, another source of static power consumption in MZMs. As a result, the MZM can be driven like a lumped element device with no need for DC routes intended to supply doped heaters. It may also be helpful when numerous devices must coexist in a limited footprint, posing the risk of thermal crosstalk during individual device operations for conventional MZMs. In the case of EO polymers, a scalable method for introducing local heat that does not de-pole neighbouring devices has yet to be established. The poling-free nature of FN-LC mitigates this issue by eliminating the necessity for a locally raised energy level. Each device can be activated using dedicated RF lines routed for normal functioning, which is of special interest for addressing PICs in many applications (e.g., cryogenic), including those with a limited number of total RF and DC I/O lines.
Towards a truly integrated SiP-FN-LC chip on a large scale, one would require to show the cooperation of several EOPs in relevant scenarios such as the one reported in [57], which require a comprehensive electrical packaging scheme and will be covered in future work. Previously, we have also successfully integrated other photonic components, such as distributed feedback laser [58] and semiconductor laser amplifier [59], using a similar photonic packaging technique shown in this work. Furthermore, we anticipate some sealing would be beneficial to accomplishing a properly packaged FN-LC SiP chip despite our early observations of the insensitivity of FN-LC to agents such as humidity and oxygen.